Re: Shift report 20000723 08:00-16:00

From: Konstantin Olchanski (olchansk@ux1.phy.bnl.gov)
Date: Sun Jul 23 2000 - 20:16:43 EDT

  • Next message: Konstantin Olchanski: "RIP http://opus.brahms.bnl.gov"

    On Sun, Jul 23, 2000 at 04:39:01PM -0400, http account wrote:
    > 
    > ------------------------------------------------------------
    > Date: Sun Jul 23 16:39:00 2000, Shift: 08:00-16:00
    > Supervisor: FV, with: DO,ZIS
    > Log book#3 pages 52 - 59
    > 
    > Shift Summary:
    > ------------------------------------------------------------
    > 2.Few achivements:
    > - BRAHVO still has a problem with H1,C2,H2,TOFW mainframe. 
    
    
    The HV controls problem is caused by a catastrophic failure of the
    terminal server on the MRS platform ("tsir"). This terminal
    server provides access to all the L1440 HV mailframes (H1, H2, C1)
    and the L1458 mainframes (TPC and RICH).
    
    In short, "tsir" refuses all incoming connections, so brahvo cannot
    connect to the mainframes. However, unexplainably, both L1458 connections
    are still working, so we still have control over the TPC high voltages.
    
    The terminal server also refuses connections to the management port,
    so it is impossible to login and reboot it from "within". Attempts
    to cycle the power on the terminal server using the magic Radio Shack
    device failed, too (the power does not go out).
    
    The next step is to either cycle the power on the MRS platform (using
    the big circuit breaker in the FEH) or do a controlled access to the IR
    to directly cycle the power on the terminal server (and also
    verify that the Radio Shack widget is configured correctly).
    
    
    -- 
    Konstantin Olchanski
    Physics Department, Brookhaven National Laboratory, Long Island, New York
    olchansk@bnl.gov
    



    This archive was generated by hypermail 2b29 : Sun Jul 23 2000 - 20:27:46 EDT